# BALLAST: Bayesian Active Learning with Look-ahead Amendment for Sea-drifter Trajectories under Spatio-Temporal Vector Fields

### v1

@@ 5.1 Ablation Study
Figure 4: Percentage utility gap with 2 standard error bounds of Uniform, EIG, and BALLAST decisions over posterior sample number $J$ at decision times $t=3,5,7$. A percentage utility gap cut-off at $1\%$ is selected with their corresponding $J$ values indicated in text.
In Figure 4, BALLAST reached the $1\%$ utility gap before $J=20$ regardless of decision times. Also, BALLAST decisions are consistently better than the uniform and EIG decisions for almost all choices of $J$, with EIG worse than Uniform – aligning with our observation in Figure 2. Full details of this ablation study and additional ablations can be found in Section G.
@@ 5.2 Temporal Helmholtz Ground Truth
Here, the ground truth vector field is drawn from a temporal Helmholtz GP described in Section 2.2 where the temporal kernel is a Matérn $3/2$ with lengthscale $2.5$ and variance $1.0$, and the Helmholtz kernel uses two RBF kernels with variance $0.5$ and lengthscales $0.8$ and $0.5$ for potential and stream kernel respectively. The ground truth vector field is considered on a time grid $[0,10]$ with time step $0.01$ and a spatial grid of size $25\times 25$ evenly-spread on $2,2]\times[-2,2]$.
All policies are initialised uniformly at time zero, and 19 further observers are deployed every $0.5$ unit of time afterwards. The performance at each deployment is measured by the average L2 error of the vectors of the posterior predictive mean field over the spatial grid and the full set of deployment times.
For the experiment, 100 runs with 10 different sampled ground truth vector fields and 10 independent runs each are conducted. We compare the policies using the average policy rank and the iso-performance with the uniform policy as the benchmark. The policy rank considers the ranks of policies (one being the best) for each iteration, and the iso-performance considers the additional (positive or negative) number of observers needed to reach the same level of performance, averaged over each iteration’s results.
@@ 5.3 SUNTANS Ground Truth
The SUNTANS model is a high-fidelity numerical fluid mechanics model for non-hydrostatic flows (Fringer et al., 2006) and internal waves (Walter et al., 2012). Here, we use a (spatial and temporal) portion of the simulated, open-sourced vector fields from Rayson et al. (2021) as the ground truth vector field – see Figure 5 for an illustration.
The surrogate model continues to be the temporal Helmholtz GP. We set the “true” model with Matérn $3/2$ temporal kernel of lengthscale 1 and variance 15, and the spatial Helmholtz kernel with RBF potential kernel of lengthscale 5 and variance 20 and RBF stream kernel of lengthscale 4 and variance 0.01. Those hyperparameter choices are learned using subsampled observations from the ground truth. This model is also used for the BALLAST-true policy.
@@
The result in Figure 7 indicates that BALLAST with true or optimised hyperparameters consistently outperforms all other policies with noticeable margins. At the end of the deployment, the two BALLAST policies save around 2 drifters against the uniform benchmark, which yields about $22\%$ deployment cost saving.
The authors of Chen et al. (2024b) justified the two criteria in Section 4.2 by comparing them to the expected information gain policy. Therefore, it is not surprising to observe DIST-SEP performing worse than EIG both here and in Section 5.2.
@@ 6 Conclusion
We apply active learning to the Lagrangian observer deployment for learning a time-dependent vector field. After noticing the inadequacy of a direct application using EIG due to its ignorance of observation structure, we introduce BALLAST to sample hypothesised vector fields and simulate potential observation trajectories for more accurate utility measurement. A novel usage of the SPDE-GP was also developed to speed up BALLAST, which could be of independent interest. Finally, our numerical experiments provide promising results on the effectiveness of our proposed method.
One direction of extension is to employ other surrogate models. Within the Gaussian process model class, there exist other physics-informed models (Hamelijnck et al., 2021, 2024; Xu and Pan, 2024). Also, model misspecification can be further addressed using post-Bayesian ideas such as Laplante et al. (2025) and Shen et al. (2025). Additionally, one could also consider deep adaptive designs (Foster et al., 2021; Iqbal et al., 2024) to amortize the acquisition optimisation for faster decisions at deployment.
@@ Acknowledgements
RZ is supported by EPSRC-funded STOR-i Center for Doctoral Training (grant no. EP/S022252/1). RZ, LA and EC are supported by the ARC ITRH for Transforming energy Infrastructure through Digital Engineering (TIDE), Grant No. IH200100009. RZ thanks Ben Lowery for the help with accessing the compute cluster, as well as William Laplante and Adrien Corenflos for discussions on SPDE-GP.
@@ Appendix A Full BALLAST Algorithm
1:Spatial grid $R$. Terminal time $T$. Stepsize $\delta_{t}$. Temporal Helmholtz GP $f$ with kernel hyperparameter $\theta$ and its extension $\bm{f}=[f,\partial_{t}f]^{T}$. Deployment number $M$. BALLAST sample number $J$. ODE solver of choice (e.g. Euler’s method).
@@ B.1 Gaussian Process
where $\sqrt{K_{**}}$ is the matrix square root of $K_{**}$, which could be obtained using multiple methods (e.g. eigendecomposition, Cholesky decomposition) (Trefethen and Bau III, 1997). The computational cost of drawing a sample is therefore $O(N_{\text{test}}^{3})$.
@@
Using the vanilla GP formulation presented above, the computational cost of likelihood training is $O(N_{\text{obs}}^{3})$, while the cost of prediction at $N_{\text{test}}$ test points is $O(N_{\text{obs}}^{3}+N_{\text{obs}}^{2}N_{\text{test}}+N_{\text{obs}}N_{\text{test}}^{2})$.
@@ Appendix D Proof of Proposition 1
Consider the sequential experimental design problem with existing observation $\mathcal{D}$, measurement set $X$, and utility $U$ where the observations are made by Lagrangian observers (see Section H.1 for details). At any decision time $t$, the deployment position $x^{S}$ following standard utility is suboptimal w.r.t. to the true utility considering all potential observations made by the placed observer.
@@ F.4 Computational Costs for Posterior Sampling
$O((N_{\text{obs-loc}}+N_{\text{space}})^{3}N_{\text{obs-time}}).$
Subsequently, the computational cost of sampling from these posterior predictive for $N_{\text{sampleT}}$ prediction times at $N_{\text{space}}$ prediction locations would be of $O((2N_{\text{space}})^{3}N_{\text{sampleT}})$.
@@ F.5 Connection to Spatial SPDE-GP
Appendix G Ablation Studies
@@
G.1 Sample Number
This study investigates the choice of posterior sample number $J$ of the BALLAST utility. In particular, we have the following acquisition function of the BALLAST-EIG policy:
@@ G.1 Sample Number
G.1.1 Synthetic Ground Truth
@@ G.1.2 SUNTANS
G.2 Time Horizon Length
@@
The general setup is identical to that of Section G.1, and we only consider the Monte Carlo gap at decision time $t=5$, without the loss of generality. In addition to the Uniform, EIG, and BALLAST decisions, we also consider BALLAST decisions with end time $T_{\text{end}}<T$, in particular, we have $T_{\text{end}}=0.1,0.5,1,2,3$. The study is conducted using 100 different randomly generated ground truth vector fields to provide uncertainty quantification of the utility gaps.
@@
Appendix H Additional Experiment Details
@@
H.1 Observations from Lagrangian Observers
@@ H.3 Considered Policies
The BALLAST-true policy implements Algorithm 1 where the GP hyperparameters are not optimised (i.e. Step 4 is skipped) but uses pre-determined, true values. The sample number $J$ is set to $20$ following the ablation results in Sections 5.1 and G.1. The projection horizon is set to be the terminal time $T$ following the ablation result in Section G.2.

### v2

@@ 5.1 Ablation Study
Figure 4: Percentage utility gap with 2 standard error bounds of Uniform, EIG, and BALLAST over posterior sample number $J$ at decision times $t=3,5,7$. A percentage utility gap cut-off at $1\%$ is selected with corresponding $J$ values in text.
In Figure 4, BALLAST reached the $1\%$ utility gap before $J=20$ consistently. Also, BALLAST decisions are consistently better than the uniform and EIG decisions for almost all choices of $J$, with EIG worse than Uniform – aligning with our observation in Figure 2. Full details of this ablation study and additional ablations can be found in Section H.
@@ 5.2 Temporal Helmholtz Ground Truth
Here, the ground truth vector field is drawn from a temporal Helmholtz GP described in Section 2.1 where the temporal kernel is a Matérn $3/2$ with lengthscale $2.5$ and variance $1.0$, and the Helmholtz kernel with two RBF kernels of variance $0.5$ and lengthscales $0.8$ and $0.5$ for potential and stream kernel respectively. The ground truth field is considered on a time grid $[0,10]$ with time step $0.01$ and a spatial grid of size $25\times 25$ evenly-spread on $2,2]\times[-2,2]$.
All policies (expect SOBOL) are initialised uniformly at time zero, and 19 further observers are deployed every $0.5$ unit of time afterwards. The performance is measured by the average L2 error of the vectors of the posterior predictive mean field over the spatial grid and the full set of deployment times. For the experiment, 10 different sampled ground truth vector fields with 10 independent runs each are conducted. We compare the policies using the average policy rank and the iso-performance with the uniform policy as the benchmark. The policy rank considers the ranks of policies (one being the best) for each iteration, and the iso-performance considers the additional (positive or negative) number of observers needed to reach the same level of performance, averaged over each iteration’s results.
@@ 5.3 SUNTANS Ground Truth
The SUNTANS model is a high-fidelity numerical fluid mechanics model for non-hydrostatic flows (Fringer et al., 2006) and internal waves (Walter et al., 2012). Here, we use a (spatial and temporal) portion of the simulated, open-sourced vector fields from Rayson et al. (2021) as the ground truth vector field – see Figure 5 for an illustration. A temporal Helmholtz GP continues to be the surrogate model of choice here for the active learning policies.
For EIG and BALLAST-true policies in particular, as they require pre-specified hyperparameter values for the GP, we set the temporal kernel to be of lengthscale 1 and variance 15, the potential component of the Helmholtz kernel to be an RBF with lengthscale 5 and variance 20, and the stream component to be an RBF with lengthscale 4 and variance 0.01. Those hyperparameter choices are selected based on estimations using ground truth observations.
@@
The result in Figure 7 indicates that BALLAST-true and BALLAST-opt consistently outperforms other policies with noticeable margins. At the end of the deployment, the two BALLAST policies save around 2 drifters against the benchmark, which yields about $22\%$ cost saving. We also notice DIST-SEP performing worse than EIG both here and above. As this heuristic was proposed as an approximation to information gain (see Section 4.2 of Chen et al. (2024b)), we find DIST-SEP’s underperformance unsurprising.
@@ 6 Conclusion
We apply active learning to the Lagrangian observer deployment for spatio-temporal vector fields. After noticing the inadequacy of standard methods, we introduce BALLAST to sample hypothesised vector fields and simulate potential observation trajectories for better utility measurement. We also propose VaSE to speed up computation, which could be of independent interest. Finally, our numerical experiments provide promising results on our method’s effectiveness.
General Applicability
While BALLAST is motivated by active learning with drifters, the proposed method can be extended to other scenarios. The deployment of collar sensors for animal movement (Handcock et al., 2009) and balloon sensors for meteorological data (Wang et al., 2020) are examples where a BALLAST-style policy can be applied. Our proposed method also contributes to the growing literature of sequential designs with search constraints (Folch et al., 2024; Mutny et al., 2023; Qing et al., 2024).
Possible Extensions
One extension is to employ other surrogate models. Within the Gaussian process model class, there exist other physics-informed models (Hamelijnck et al., 2021, 2024; Xu and Pan, 2024). Additionally, one could also consider deep adaptive designs (Foster et al., 2021; Huang et al., 2024; Iqbal et al., 2024) to amortize the acquisition optimisation for faster decisions at deployment.
Impact Statement
This paper presents work whose goal is to advance the field of machine learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ Acknowledgements
RZ is supported by EPSRC-funded STOR-i Center for Doctoral Training (grant no. EP/S022252/1). RZ, LA and EC are supported by the ARC ITRH for Transforming energy Infrastructure through Digital Engineering (TIDE), Grant No. IH200100009. RZ thanks Ben Lowery for the help with accessing the compute cluster, as well as William Laplante, Adrien Corenflos, and Matthias Sachs for discussions on SPDE-GP.
@@ Appendix A Full BALLAST Algorithm
1:Input: Spatial grid $R$. Terminal time $T$. Stepsize $\delta_{t}$. Temporal Helmholtz GP $f$ with kernel hyperparameter $\theta$ and its extension $\bm{f}=[f,\partial_{t}f]^{T}$. Deployment number $M$. BALLAST sample number $J$. ODE solver of choice (e.g. Euler’s method).
@@ B.1 Gaussian Process
where $\sqrt{K_{**}}$ is the matrix square root of $K_{**}$, which could be obtained using multiple methods (e.g. eigendecomposition, Cholesky decomposition) (Trefethen and Bau III, 1997).
@@
Using the vanilla GP formulation presented above, the computational cost of likelihood training is $O(N_{\text{obs}}^{3})$, while the cost of prediction at $N_{\text{test}}$ test points is $O(N_{\text{obs}}^{3}+N_{\text{test}}^{3}+N_{\text{obs}}^{2}N_{\text{test}}+N_{\text{obs}}N_{\text{test}}^{2})$.
@@ Appendix D Proof of Proposition 1
Consider the sequential experimental design problem with existing observation $\mathcal{D}$, measurement set $X$, and utility $U$ where the observations are made by Lagrangian observers (see Section I.1 for details). At any decision time $t$, the deployment position $x^{S}$ following standard utility is suboptimal w.r.t. to the true utility considering all potential observations made by the placed observer.
@@ F.4 Computational Costs for Posterior Sampling
$O((N_{\text{obs-loc}}+N_{\text{space}})^{2}N_{\text{obs-time}}).$
Subsequently, the computational cost of sampling from these posterior predictive for $N_{\text{sampleT}}$ prediction times at $N_{\text{space}}$ prediction locations would be of $O((2N_{\text{space}})^{2}N_{\text{sampleT}})$.
@@ F.5 Connection to Spatial SPDE-GP
Appendix G Vanilla SPDE Exchange Details
Let $f$ be a spatio-temporal GP with separable kernel $k=k_{s}k_{t}$ where $k_{s}$ is the spatial kernel and the time kernel $k_{t}$ is Matérn (with smoothness $3/2$ here, but can be extended to any half-integer smoothness). We are interested in a spatial grid $R$ and temporal grid $\mathcal{T}$ that is a regularly interval timestamps in $[0,T]$ with time gap $\delta$. The temporal grid does not have to be regular, but for the simplicity of presentation this is assumed. A non-regular temporal grid would only require us keeping track of the time increments and adjusting the SPDE-GP updates accordingly, which do not affect the general vanilla SPDE exchange (VaSE) method.
Consider at time $t_{n}\in\mathcal{T}$ we have observations $\mathcal{D}$. The VaSE method to sample from the posterior $f|\mathcal{D}$ at full spatio-temporal grid for time $t_{n}$ onwards consists of three parts:
  1. 1.
    posterior GP update with $\mathcal{D}$ for the extended GP $\bm{f}=[f,f^{\prime}]^{T}$ using standard GP,
  2. 2.
    compute the posterior predictive distribution of $\bm{f}$ at $R\times\{t_{n}\}$,
  3. 3.
    sample from the posterior predictive and propagate for the remaining timestamps using SPDE-GP.
The first and last parts apply the vanilla and the SPDE methods, while the middle part acts as the exchange to switch between the two modes of GP inferences. Details of the SPDE-GP was presented in Section F.
When the smoothness parameter of the temporal Matérn kernel becomes $(2j+1)/2$ for some nonnegative integer $j$, we will adjust the above procedure with an extended GP $\bm{f}=[f,f^{\prime},\cdots,f^{(j)}]^{T}$ with all $j$ derivatives. This is to ensure the SPDE-GP propagation has the necessary dimensions.
G.1 Computational Cost Analysis and Comparison
We conduct a cost analysis here. Let $N_{\text{obs}}$ denotes the number of observations, $N_{\text{obs,t}}$ denotes the number of distinct observation time slices, $N_{\text{pred,s}}$ and $N_{\text{pred,t}}$ denotes the number of prediction spatial and temporal points. Below, we will carefully present the computational cost of drawing a posterior sample under the standard GP, SPDE-GP, and VaSE where the cost orders for each portion of the sampling cost are included in square brackets.
Following from Section B.1, the standard GP posterior sample requires computing the posterior predictive distribution’s covariance $N_{\text{obs}}^{3}$] and taking its matrix square root $N_{\text{pred,s}}^{3}N_{\text{pred,t}}^{3}$], as well as multiplying the Gaussian noise $N_{\text{obs}}^{2}N_{\text{pred,s}}N_{\text{pred,t}}$] and adding the computed posterior mean $N_{\text{obs}}N_{\text{pred,s}}^{2}N_{\text{pred,t}}^{2}$].
Following from Section F, the SPDE-GP approach to draw a posterior sample consists of first creating update matrices via Kronecker products of the temporal update matrices and the full Gram matrix of all $N_{\text{pred,s}}+N_{\text{obs}}$ considered spatial points, followed by matrix inversions for assimilating observations $(N_{\text{obs}}+N_{\text{pred,s}})^{3}N_{\text{obs,t}}$], then using the coordinates corresponding to test points to scale Gaussian noise and propagate forward to generate the sample $N_{\text{pred,s}}^{2}N_{\text{pred,t}}$].
Finally, the VaSE approach generates the mean vector and covariance matrix of the posterior predictive distribution at the current time over all spatial test points $N_{\text{obs}}^{3}+N_{\text{pred,s}}^{2}N_{\text{obs}}+N_{\text{pred,s}}N_{\text{obs}}^{2}$], then this spatial component is used to scale Gaussian noise and propagate forward to generate the sample $N_{\text{pred,s}}^{2}N_{\text{pred,t}}$].
In summary, the three methods’ costs to generate one posterior sample are
$\begin{split}&(\textbf{Standard})~O(N_{\text{obs}}^{3}+N_{\text{pred,s}}^{3}N_{\text{pred,t}}^{3}+N_{\text{obs}}^{2}N_{\text{pred,s}}N_{\text{pred,t}}\\ &\qquad\qquad\quad+N_{\text{obs}}N_{\text{pred,s}}^{2}N_{\text{pred,t}}^{2}),\\ &(\textbf{SPDE})~O((N_{\text{obs}}+N_{\text{pred,s}})^{3}N_{\text{obs,t}}+N_{\text{pred,s}}^{2}N_{\text{pred,t}}),\\ &(\textbf{VaSE})~O(N_{\text{obs}}^{3}+N_{\text{pred,s}}^{2}N_{\text{obs}}+N_{\text{pred,s}}N_{\text{obs}}^{2}\\ &\qquad\qquad+N_{\text{pred,s}}^{2}N_{\text{pred,t}}).\end{split}$
Appendix H Ablation Studies
@@
H.1 Sample Number
This study investigates the choice of posterior sample number $J$ of the BALLAST utility. In particular, we have the following acquisition function of BALLAST-EIG:
@@ H.1 Sample Number
H.1.1 Synthetic Ground Truth
@@ H.1.2 SUNTANS
H.2 Time Horizon Length
@@
The general setup is identical to that of Section H.1, and we only consider the Monte Carlo gap at decision time $t=5$, without the loss of generality. In addition to the Uniform, EIG, and BALLAST decisions, we also consider BALLAST decisions with end time $T_{\text{end}}<T$, in particular, we have $T_{\text{end}}=0.1,0.5,1,2,3$. The study is conducted using 100 different randomly generated ground truth vector fields to provide uncertainty quantification of the utility gaps.
@@
Appendix I Additional Experiment Details
@@
I.1 Observations from Lagrangian Observers
@@ I.3 Considered Policies
The BALLAST-true policy implements Algorithm 1 where the GP hyperparameters are not optimised (i.e. Step 4 is skipped) but uses pre-determined, true values. The sample number $J$ is set to $20$ following the ablation results in Sections 5.1 and H.1. The projection horizon is set to be the terminal time $T$ following the ablation result in Section H.2.